### v1

# First-Principles Study of the Fermi Surface Topology of CeCu2Si2

Abstract
Since the discovery of heavy-fermion superconductivity in CeCu2Si2, the material has attracted great interest particularly with regard to the nature of the superconducting pairing and its mechanism. Consequently, it is essential to better understand the electronic Fermi surface topology and its role in strong antiferromagnetic fluctuations. The standard density functional theory method is insufficient to model the interplay of strong onsite Coulomb repulsion in localized 4f-electrons and their hybridization with itinerant ligand-orbital electrons. We have performed electronic ground state calculations on CeCu2Si2 using the Gutzwiller wavefunction approximation. The Gutzwiller approximation captures the quasiparticle band renormalization from the strong onsite Coulomb repulsion. We have performed an analysis of this effect on the electronic structure and the Fermi surface topology by varying the interaction strength and taking into account the crystal-field splitting. Using the de Haas van Alphen effect, the extremal Fermi surface cross-sectional areas were calculated to quantify the effects of quasiparticle mass renormalization on the Fermi surface. Our results confirm two Fermi surface sheets corresponding to the heavy (488me) and light (4.35me) quasiparticles, which is in close agreement with experimental measurements as well as the renormalized band method.
1 Introduction
The unconventional superconductor (SC), CeCu2Si2, has been a material of interest for research in both the SC community and the heavy fermion community for over forty-five years. In 1979 Steglich et al. published their findings on CeCu2Si2 showing a low temperature transition into a superconducting state at around 0.5 K [1]. This marked the first demonstration of superconducting pairing occurring in a metal driven by many-body interactions. It was postulated in this discovery work that the interactions would be “probably magnetic in origin" arising from the low temperature anomalies characterized by an instability of the 4f shell. Subsequent studies on CeCu2Si2 would point to the importance of the Fermi surface and crystal field splitting in understanding the nature of the f electron behavior [2, 3, 4].
In 1990, magnetic oscillations were observed for CeCu2Si2, and the de Haas Van Alphen (dHvA) effect was measured for one of the quasiparticle sheets making up the Fermi surface. They approximated quasiparticles with an effective mass of 4.5-6 times the mass of an electron[5]. The measurement of the Fermi surface through dHvA effects is one of most useful ways to understand the f electrons directly. Therefore, shortly after these measurements were published, Hirama and Yanase presented density function theory (DFT) calculations in the local density approximation (LDA) [6, 7], which predicted a Fermi surface comprised of three sheets. Though they were unable to conclusively determine how their calculations compared with experiments at the time, their work was foundational to the understanding of electronic states in CeCu2Si2. A year later, Zwicknagl et al.[8] published an electronic structure study using renormalized band theory. This method predicted two sheets in the Fermi surface,one with light quasiparticles and the other with heavy quasiparticles. The light quasiparticle sheet was in good agreement with the previous LDA calculations, but the heavy sheet was found to be highly sensitive to the renormalizing effects of strongly correlated electrons. Their study also suggested that the superconducting phase transition is likely driven by a Fermi surface topological transition in the heavy quasiparticle system.
Many of these early predictions started to be confirmed with improved sample preparation techniques and theoretical methods [9, 10, 11, 12, 13]. In 2011, neutron scattering experiments identified antiferromagnetic excitations as the driving force for superconducting pairing in CeCu2Si2 [14]. Unlike conventional SCs where superconductivity is destroyed by magnetic ions, CeCu2Si2 relies on a periodic dense array of magnetic moments from the Ce3+ ions’ electrons to transition to a superconducting state [15]. Given the confirmed importance of the electrons in the physics of the system, in 2014, dynamical mean field theory (DMFT) was applied to CeCu2Si2 to attempt a better explanation of the heavy fermion physics of the material and explore the importance of the crystal field splitting [16]. This method predicted a single heavy fermion sheet in the Fermi surface with a quasiparticle effective mass of 200me. Zwicknagl et al.[17] once again approached the material with renormalized band theory in 2016, predicting, again, a Fermi surface comprised of two sheets, with a light quasiparticle effective mass of 5me and a heavy quasiparticle mass of 500me.
Several theoretical studies have confirmed the importance of the electron correlations and crystal field effects in CeCu2Si2[9, 12, 18, 19, 20, 21], including the work of Amorese et al. in 2020[22], which does a through study of the crystal field states using renormalized band theory and X-ray absorption spectroscopy (XAS). Subsequently, Li et al.[23] applied LDA+U within the random phase approximation (RPA)[24] to predict two Fermi surface sheets in good agreement with Zwicknagl et al.[8, 17], and Li et al.[20] and Luo et al.[25] applied DMFT to the electronic analysis and temperature dependence respectively. Additional works have been done exploring the quasiparticle behavior and impurity effects using DFT+[26, 27], though these calculations do not take into account the crystal field splitting. The existence of the heavy quasiparticle Fermi surface sheet has been experimentally observed as of 2021 in Wu et al.[28] using angle resolved photo emission spectroscopy (ARPES), which qualitatively confirms the presence of the two Fermi surface sheets calculated by both renormalized band theory and LDA+U methods, though in this case their calculations predicted the mass renormalization of the heavy band to be .
Here we present an alternative theoretical approach to calculate the Fermi surface topology and mass enhancement for CeCu2Si2 using the Gutzwiller Wavefunction Approximation (GWA)[29]. Through this investigation we show that the GWA is capable of reproducing the two Fermi Sheets indicated by experiment and that crystal field splitting (CFS) is as significant to the underlying physics of the electronic structure of CeCu2Si2 as the electronic correlations and spin-orbit coupling (SOC), which we demonstrate through comparisons to GWA calculations that do not take CFS into account. We have also determined the sign of the charge and crystal field parameter, , to be in agreement with the X-ray scattering measurements of Willers et al.[30] and Rueff et al.[31]. Furthermore, the increased computational efficiency of the GWA method allows us to perform calculations on complex systems, such as CeCu2Si2, with less computational expense than DMFT [32, 33, 34]. In the following sections, we describe the theoretical framework for addressing the CFS within the GWA (section 2.1), the results of our electronic structure calculations and analysis (section 2.2), and our analysis of the Fermi surface topology (section 2.3). We compare our results for calculations performed with and without electronic correlations and with and without CFS incorporated into the correlation effects. A conclusion is given in section 3. The details of our methodology are described in section 4. A full account of our analysis of the Coulomb parameter tests is given in the Sumplementary Material[35].
2 Results and Analysis
2.1 Theory: Crystal Field Splitting
The atomic environment of the CeCu2Si2 structure is analyzed by taking into account the crystal field splitting for the Ce atom in a body centered tetragonal (BCT) environment of 8 Si atoms approximated as point charges (FIG. 1). These neighboring charges contribute to an additional potential, , on the Ce atom. When an atom (or ion) is surrounded by point charges at location , an electron of charge at location of this atom is subject to an additional electrostatic potential given by
where is the charge from the atomic species around Ce (i.e. Si in the first shell of neighboring charges and Cu in the second). The surrounding silicons have wave functions spreading beyond their atomic site, making it difficult to obtain a precise estimate of . We only include the first shell of neighboring charges, whereas the further shells with alternating charge signs could also modify the amplitude of the splitting (but not the general symmetry). We have tested these alternatives to verify that the overall sign of the charge does not change[35].
All charges of silicon are identical and are at the same distance of the central atom. From ab initio DFT calculations (see section 4) we have determined that [36]. Our results are consistent with those given by group theory[37], and a full derivation is given in section I.A of the Supplementary Materials[35]. To summarize, the ground state of the Ce atom can be split into three Kramers doublets, two and a as shown in FIG. 1(c). Since is positive we can deduce that , which is consistent with the findings of Willers et al.[30] and Rueff et al.[31]. Thus the ground state doublet can be written in the basis as
Then the doublet
gives the first excited states, with the energy gap between this doublet and the ground state doublet, experimentally known to be of the order of 30 meV. Finally, the highest energy states constitute the doublet
For the electronic structure calculations discussed in the following sections, the CFS effects were calculated using the implementation in the Cygutz code as detailed in Lanatà et al[38].
2.2 Electronic Structure
The electronic band dispersion, electron density of states (DOS), and Fermi surface for CeCu2Si2 were calculated within the GWA with and without electronic correlations. The calculations use a DFT basis as described in section 4. Since DFT includes CFS, all of the initial calculations take the crystal field effects into account at the weakly correlated level. To better understand the interplay between the electron correlation and the CFS, we have also tested the GWA with and without CFS included in the band renormalization parts of the calculation[38]. The results without correlations () or additional CFS are given in FIG.2 (a) and (d). The three corresponding Fermi surface sheets are given in FIG.3 in order of descending band number from top to bottom for each doubly degenerate band. These results are in close agreement with those calculated in the early 1990’s by Harima et al[6, 7]. The results with electronic correlations and CFS are shown in FIG.2 (b) and (e) with a Coulomb interaction strength of eV and in (c) and (f) with eV. An exchange coupling parameter of eV was used for each calculation where and relativistic (SOC) effects are considered in a paramagnetic configuration for each case (see section 4.1).
The localization of the f-electrons can be seen both in the band dispersion and the DOS. For the case without electron-electron correlations or additional CFS, the two spikes in the DOS just above the Fermi energy are the and orbital sub-shells, respectively (FIG.2d). This indicates that the orbital character at the Fermi energy is dominated by the six-fold degenerate state of the 4f-electron. When electronic correlations ( eV) are taken into account, the CFS effect is enhanced. As such, the DOS has a sharp peak at the Fermi energy and a more disperse arrangement of the higher energy quasiparticle states (FIG.2e), which is consistent with the lifting of the Ce electrons due to CFS (FIG. 1). This appears in the band dispersion (FIG.2b) as a broadening of the f-electron occupation and a flattening of the bands around the Fermi energy leading to two, rather than three, sheets in the Fermi surface (FIG.3b). Calculations for eV and eV are shown in FIG.2 to demonstrate the sensitivity of the electronic ground state to the presence and strength of the electron-electron interactions. The effects of CFS can be more clearly seen in the analysis of the Fermi surface in FIG.3, which is discussed in section 2.3.
As the electron correlations become stronger in the system, the peak grows sharper and narrower, becoming more delta-like. Conversely, the occupation appears at a higher energy while becoming more disperse as increases to 6 eV. This is consistent with expected SOC splitting in a strongly coupled system. A full analysis of the Coulomb parameter strength was performed for cases to eV and is discussed in more detail in section III.A of the Supplementary Material[35], but the impact of the electronic correlations are most dramatic between eV and eV where the system itself begins to have precarious stability eventually leading to a collapse of the bands around the Fermi energy once the correlation strength exceeds a critical point between eV. For our particular method, we found a critical interaction strength of eV gave the best overall convergence and agreement with past studies [8, 17, 22, 28, 19, 27].
Notably, the sensitivity of the electronic structure is even more pronounced in the presence of CFS as shown by the eigenvalues of the 4f quasiparticle matrix, , given in Table 1 along with the corresponding orbital occupations, . is the matrix of quasiparticle weights derived from the "Gutzwiller self-energy" defined by equation 14 of Lanatà et al.[38] (and discussed in section III.A.1 of the Supplementary Material[35]) to describe the rotationally invariant slave boson (RISB) mean-field theory[32, 38]. The eigenvalues of this matrix can be related to the effective mass of the quasi particle as . These eigenvalues have been calculated with and without CFS effects and recorded in Table 1, which compares these quasiparticle weights in each case for increasing values of . In the case where CFS is neglected, the f electrons are split into and sub-orbitals, and the magnitude of decreases as increases indicating a gradual increase in the effective quasiparticle mass. The orbital occupation increases very slightly with increasing in the subshell and decreases with increasing in the subshell. For the calculations including CFS, the subshell is split into the , , and states (FIG. 1). In the Kramers doublet, defined by equation 2, decreases sharply between eV and eV. Here we see that this slight increase of the Coulomb interaction strength corresponds to a relatively sharp decrease in the quasiparticle weight in and corresponding increase in the effective quasiparticle mass, . This is consistent with the evolution of the DOS and our exploration of the Fermi surface topology (see section 2.3), which indicates an increasingly sensitive and coupled relationship between the electronic correlations and the CFS between eV and 6 eV. Indeed, at eV, approaches zero. The total orbital occupation is consistent with a localized 4 configuration with the greatest occupancy in the doublet. This is in agreement with XPS[39] and XAS[40] observations as well as previous theoretical analysis based on DMFT calculations[20]. An account of the complete results for the quasiparticle weights and orbital occupations from eV to 6 eV is given in the Supplementary Material[35].
| With CFS | ||||||
| U (eV) | Z | n | Z | n | Z | n |
| 0.0 | 0.96 | 0.48 | 0.96 | 0.02 | 0.96 | 0.04 |
| 5.1 | 0.47 | 0.50 | 0.96 | 0.01 | 0.99 | 0.01 |
| 6.0 | 0.00 | 0.50 | 0.97 | 0.01 | 0.96 | 0.01 |
| Without CFS | ||||||
| U (eV) | Z | n | Z | n | ||
| 0.0 | 0.96 | 0.10 | 0.96 | 0.04 | ||
| 5.1 | 0.48 | 0.14 | 0.74 | 0.02 | ||
| 6.0 | 0.32 | 0.15 | 0.75 | 0.01 | ||
2.3 Fermi Surface Topology
Analysis of the Fermi surface topology gives a clear indication of the impact of the CFS in CeCu2Si2. There are numerous theoretical studies and experimental measurements that look at the Fermi surface. Based on the combined efforts of magnetic oscillation studies [5] and ARPES experiments[28], it has been experimentally confirmed that there are two sheets in the Fermi surface. Theoretical methodologies from the LDA[6, 7] to LDA+ in the random phase approximation (RPA)[24, 18, 19, 28, 27] to renormalized band theory[8, 17] and DMFT[16] have been applied to understand the Fermi surface topology. The theoretical results appear to depend heavily on the method, choice of electronic correlation strength, and the presence of CFS.
In the early works of Harima et al.[6, 7], the LDA method predicted three Fermi surface sheets. Our GGA calculations at , agree closely with both the qualitative character of the Fermi sheet topologies as well as the dHvA simulation results for the maximal frequencies and cyclotron masses. In our survey of the Hubbard parameter strength in the GWA[35], we have tracked the evolution of the Fermi surface topology (FIG. 3), and simulated the dHvA extremal frequencies for an external B field parallel to the -axis. The dHvA results for the Fermi surface obtained using the simulation method of Rourke and Julian[41] described in section 4.2 are given in Table 2, which shows the results when CFS is included and when it is neglected. For , the impact of CFS on the Fermi surface topology is negligible.
It is well documented that better agreement with the current experimental consensus can be achieved in the Fermi Surface by including strong electron-electron correlations (via a Coulomb interaction). The DFT++RPA type methods tend to over estimate the impact of the 4f electrons in the conduction bands, but they are capable of giving results that compare qualitatively well with experiment. These methods also appear to be successful with smaller values according to Ikeda et al.[18]. In our calculations using a GWA approach without CFS the correlation strength needs to be large in order to predict a Fermi surface containing two sheets, rather than three. Even at eV this method does not give the expected cylindrical heavy band predicted by the DFT++RPA and renormalized band theory methods (FIG. 3). Additionally, we explored a DMFT approach (neglecting CFS) discussed in the Supplementary Materials[35]. These Fermi surface topology results compare nearly exactly to our GWA calculation at the same value of , which is shown in FIG.S10. There have been several DMFT studies of CeCu2Si2[16, 20, 25]. The study by Pourovskii et al..[16] included CFS and predicted a single sheet in the Fermi surface, which is at odds with previous DFT+ methods and experimental measurement. It is, however in close agreement with our GWA+CFS results at eV (FIG. 3).
The electronic structure results from section 2.2 are an indication that the CFS effects are significant to the band dispersion and therefor the Fermi surface. This is demonstrated by the comparison of FIG.3 (a) and (b), where (b) shows the evolution of the Fermi surface topology with increasing with CFS included. At eV with CFS, there are two sheets making up the Fermi surface and the dHvA data is given in Table 2. The qualitative Fermi surface is in close agreement with the renormalized band theory studies of Zwicknagl et al.[8, 17], as well as the DFT++RPA methods[18, 19]. The effective cyclotron mass (eq. 6) calculated with CFS also agrees closely with the renormalized band theory calculations which predicted an effective mass of for the light quasiparticle sheet and for the heavy sheet. Our calculations predict 4.35 and 488 for the light and heavy sheets respectively (see section III.B for the Suplementary Material[35]).
| with CFS | without CFS | ||||||||
| U (eV) | band | e- | e- | ||||||
| 0.0 | 53 | 10.49 | 3.282 | 0.239 | 1.840 | 10.49 | 3.282 | 0.242 | 1.838 |
| 55 | 3.397 | 2.480 | 0.210 | 0.141 | 3.397 | 2.480 | 0.215 | 0.144 | |
| 57 | 0.093 | 0.884 | 0.002 | 0.001 | 0.093 | 0.884 | 0.002 | 0.001 | |
| 5.1 | 53 | 11.45 | 4.353 | 0.518 | 1.653 | 10.35 | 9.214 | 0.203 | 1.864 |
| 55 | 5.910 | 488.0 | 1.007 | 0.676 | 2.741 | 72.230 | 0.149 | 0.100 | |
| 57 | - | - | - | - | 0.357 | 3.419 | 0.005 | 0.003 | |
| 6.0 | 53 | 18.346 | 2000 | 2.127 | 1.427 | 10.85 | 15.908 | 0.232 | 1.844 |
| 55 | - | - | - | - | 3.333 | 66.60 | 0.153 | 0.103 | |
3 Conclusion
In the last years there have been a number of studies using various methods to describe the Fermi surface of CeCu2Si2. Experimental measurements have confirmed the existence of a Fermi Surface comprised of two sheets[5, 28], several theoretical and experimental studies have explored the importance of the CFS and electron correlation effects on the overall electronic structure and the superconducting pairing[9, 12, 41, 30, 18, 19, 20, 22], and new advances have made it possible to perform more accurate and detailed analysis of the intricate coupling and correlation effects in this complex material.
In this study we have presented a systematic analysis of the electronic structure and Fermi surface topology of CeCu2Si2 as a function of the Coulomb interaction strength with a focus on the CFS effect. We have performed calculations using the GWA to show that the resulting Fermi surface topology is highly sensitive to methodology, electron-electron interactions, and CFS. De Haas van Alphen simulations were applied to the Fermi surfaces to analyze the Fermi sheets. Our GWA results calculated using a Hubbard parameter of eV and including CFS are found to be in agreement with both experimental measurements as well as the renomalized band theory. More broadly, our results demonstrate that the CFS and electronic correlations are on the same order of importance in the electronic structure of CeCu2Si2. The qualitative and quantitative aspect of the Fermi surface are also extremely sensitive to small changes in the electron correlation, particularly in the presence of CFS. Based on the analysis of the electronic DOS, this is due to the complexity of the 4f electron behavior present in the heavy Fermi sheet.
Our results highlight the necessity for a thorough treatment of the electronic coupling in order to better understand the nature of the heavy-fermion superconductivity and formation of Cooper pairs. Future studies of the temperature dependence (which has been touched on in Luo et al..[25]) and possible electron-phonon coupling contribution would aid in forming an even more comprehensive picture of CeCu2Si2.
4 Methods
4.1 Computational Methods
To account for the strong electronic coupling in CeCu2Si2 we combine standard density functional theory with the GWA method implemented in the CyGutz code [32, 38]. This method uses the full-potential, linearly augmented plane wave (FPLAPW) DFT of WIEN2k [42] as its basis and implements a combination of the GWA and the Slave Boson method to account for strong electronic correlations. Relativistic spin-orbit coupling (SOC) effects were included with a k-point grid of and muffin tin radii of 2.50 for the Ce and Cu atoms and 1.93 for the Si atoms, where is the Bohr radius. The cut-off parameter was set to . A generalized gradient approximation (GGA-PBE)[43] exchange correlation functional was used throughout all calculations. The Coulomb interaction strength was tested at values from 0 eV to 6.0 eV[35]. We tested the GWA band renormalization implementation with and without CFS effects included beyond the initial DFT. For all calculations where , Hund’s exchange coupling parameter, , was set at 0.7 eV. For comparison, a GGA+DMFT calculation was performed for eV and eV. Results and details of the methodology are given in the Supplementary Materials, section II[35].
4.2 de Haas van Alphen simulations
Analysis of the Fermi surfaces was done using numerical calculations of the dHvA effect as implemented by Rourke and Julian[41]. By applying a magnetic field to the system, oscillations in the magnetic susceptibility can be determined from the changes in the number of occupied Landau levels as a function of the reciprocal magnetic field, [41, 44]. Then the dHvA frequency can be expressed as
where is the elementary charge of an electron, and is the extremal cross-sectional area of the branch of the Fermi surface in a plane perpendicular to . The effective carrier mass averaged around the extremal cyclotron orbits is also determined from
where is in units of the electron mass, . This gives the instantaneous effective mass of the quasiparticles in the Fermi sheet at a particular orbit rather than the overall effective mass of the quasiparticles in the energy band. The results of the dHvA analysis are given in Table 2.
Acknowledgements
We thank Chao Cao, Bogyu Jang, E. M. Nica, F. Steglich and G. Zwicknagl for useful discussions. This work was carried out under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration under Contract No. 89233218CNA000001. The Fermi-surface topology analysis work was partly supported by the LANL LDRD Program and the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. The GWA simulation work was supported by the NNSA Advanced Simulation and Computing Program. It was, in part, supported by the Center for Integrated Nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resource as well as NERSC.The work at Rice has primarily been supported by the National Science Foundation under Grant No. DMR-2220603, by the Robert A. Welch Foundation Grant No. C-1411 and the Vannevar Bush Faculty Fellowship ONR-VB N00014-23-1-2870. Q. S. acknowledges the hospitality and support by a Ulam Scholarship from the Center for Nonlinear Studies at Los Alamos National Laboratory during the initial stage of the work.
Author contributions statement
R.M.T performed calculations and prepared the results and analysis. J-P.J formulated the theory for the Crystal Field analysis and advised on GWA theory. Q.S formulated the initial motivation for the work and advised on conceptualization for the project. J-X.Z formulated the initial motivation for the project and advised at all levels. All authors participated in the discussion of results and reviewed the manuscript.
Additional information
The authors declare that they have no competing financial interests.